# Natural Language Generation from Visual Events: Challenges and Future Directions

### v2

@@
Abstract
The ability to use natural language to talk about visual events is at the core of human intelligence and a crucial feature of any artificial intelligence system. In recent years, a substantial body of work in visually grounded NLP has focused on describing content depicted in single images. By contrast, comparatively less attention has been devoted to exhaustively modeling scenarios in which natural language is employed to interpret and talk about events presented through videos or sequences of images. In this position paper, we argue that any NLG task dealing with sequences of images or frames is an instance of the broader, more general problem of modeling the intricate relationships between visual events unfolding over time and the features of the language used to interpret, describe, or narrate them. Therefore, solving these tasks requires models to be capable of identifying and managing such intricacies. We consider five seemingly different tasks, which we argue are compelling instances of this broader multimodal problem. Consistently, we claim that these tasks pose a common set of challenges and share similarities in terms of modeling and evaluation approaches. Building on this perspective, we identify key open questions and propose several research directions for future investigation. We claim that improving language-and-vision models’ understanding of visual events is both timely and essential, given their growing applications. Additionally, this challenge offers significant scientific insight, advancing model development through principles of human cognition and language use.
1 Introduction
@@
To date, much of the progress in visually grounded NLP has focused on language generation from a static image. Tasks such as image captioning Vinyals et al. (2015) and visual question answering Antol et al. (2015) have demonstrated the potential of multimodal models to ground linguistic expressions in visual input Baroni (2016); Beinborn, Botschen, and Gurevych (2018); Wang et al. (2024). However, the world is not static. Real-world applications such as media content creation, real-time sports commentary and news reporting, and interactive multimodal assistants frequently require reasoning about events as they unfold over time. This requires capturing key aspects of event understanding, including temporal dynamics, causality, and narrative structure, that arise from complex multimodal interaction.
@@
In this position paper, we argue for a unified perspective: that all NLP tasks dealing with visual events, whether presented through videos or ordered image sets, are instances of a broader, more general problem. At the heart of this problem lies the need to model the intricate relationship between temporally ordered visual events and the structure, content, and function of the language used to interpret, describe, or narrate them. We claim that solving these tasks effectively requires models to be capable of identifying and managing such intricacies, leveraging the temporal, causal, and discourse-level patterns inherent in both visual data and language, as well as the many ways in which these aspects interact.
@@
To ground our argument, we examine five seemingly distinct tasks, each of which we claim exemplifies this broader multimodal challenge. Our aim is not to provide a survey of relevant tasks, but rather to illustrate the general problem with a few representative tasks. The tasks we consider are Change Captioning Jhamtani and Berg-Kirkpatrick (2018), Video Question Answering (Video QA) Zeng et al. (2017); Bansal, Zhang, and Chellappa (2020), Video Captioning Yao et al. (2015), Movie Auto Audio Description (Movie Auto AD) Han et al. (2023b, a), and Visual Storytelling Huang et al. (2016). Figure 1 presents an illustration and a prototypical example for each of the five tasks, drawn from publicly available datasets. By analyzing them through a shared lens, we highlight common modeling and evaluation challenges that are often overlooked when tasks are considered in isolation. Yet, we recognize that these tasks also differ in crucial ways that stem from their distinct communicative intents and objectives. For instance, crafting a coherent and imaginative narrative in Visual Storytelling poses fundamentally different demands compared to pinpointing specific visual changes in Change Captioning. These task-specific nuances, which can be ignored or flattened when focusing solely on shared technical challenges, underscore the need for specific modeling proposals and evaluation protocols. This constitutes the second key argument of our position paper.
@@
  • We offer a comprehensive synthesis of modeling approaches, training paradigms, and evaluation methodologies used across five vision-to-language tasks involving visual events. By identifying shared pillars and methodological commonalities, we provide a unified lens that we hope will serve as a foundation for future research.
@@
  • We show that, within the problem of generating language from visual events, different tasks vary along various features based on their communicative intents. To this end, we provide a quantitative characterization of the tasks based on two dimensions—visual similarity and textual consistency—capturing the similarity of the elements in the visual input, and the consistency of the textual output.
  • We outline several promising research directions aimed at advancing models’ ability to capture the complex, time-dependent relationships between visual input and linguistic output. These include cognitively inspired representations, improved discourse modeling, and principled evaluation metrics that reflect narrative quality.
{tblr}
@@
<Write a story for this image sequence.,
@@
2 Model Architectures
Modeling approaches to text generation from visual events have evolved over time from being based on recurrent neural networks (RNN; Hochreiter and Schmidhuber, 1997) to being transformer-based Vaswani et al. (2017). More recent models directly leverage pre-trained large (vision)-language models (LLMs/VLMs), often in a zero-shot manner. In this section, we discuss this evolution and summarize the various state-of-the-art model architectures proposed for this problem. We ground our discussion in five representative tasks—listed in Figure 1—that serve to exemplify the main trends. Change Captioning requires generating a caption to describe semantic changes between a pair of related input images. This task is particularly relevant for applications involving surveillance or quality inspection of ‘before’ and ‘after’ visual data. The goal of Visual Storytelling is to generate a fictitious, coherent narrative from a temporally-ordered sequence of images or frames. Applications of this task include digital storytelling and media creation. Video Captioning concerns describing videos by summarizing them, while Video QA involves generating appropriate answers to questions about the entities and events in videos. Both tasks have a wide range of applications, such as content summarization on streaming platforms and interacting with virtual assistants. Movie Auto AD involves generating narrative descriptions to complement the dialogues and soundtracks of movies. Enhancing the accessibility of movies and automatizing dubbing are some of the applications of this task.
@@ 2 Model Architectures
2.1 Vision Encoder
The primary purpose of a vision encoder in vision-to-language tasks is to extract information from the visual input and to optimally encode it into a contextual representation that guides language generation. To achieve this, encoders in the proposed models follow multiple steps, some of which are common across language generation tasks from visual events. First, a pre-trained vision model is utilized for extracting feature representations of the raw input sequences of images or video frames. Earlier approaches used vision models based on convolutional neural networks (CNNs) such as ResNet He et al. (2016) or R3D Tran et al. (2018) that are primarily pre-trained on the object detection task using large amounts of image or video data. Most of the recent models across tasks use transformer-based vision models pre-trained for various image-only and image-text alignment objectives, e.g., CLIP-ViT-L Radford et al. (2021). We note that besides the primary input sequence of images or video frames, models proposed for some of the tasks, e.g., Movie Auto AD, utilize additional input data such as close-ups of characters in the movie clips (exemplars) Han et al. (2023a). Furthermore, the TAPM Yu et al. (2021) model proposed for the Visual Storytelling task utilizes FasterRCNN Ren et al. (2015) for extracting such local entity-level features alongside the global image-level features from ResNet.
@@ 2.1 Vision Encoder
Beyond these common components across tasks, vision-encoders may also contain additional task-specific steps for capturing the visual information in a way that suits the task’s objective better. For instance, the ViLA model Wang et al. (2025) for the Video QA task utilizes a learnable Frame-Sampler component to efficiently select a small subset of frames that are most likely to contain the relevant information needed to answer the question. Another example with a task-specific step is the MSCM+BART model Hsu et al. (2020) for Visual Storytelling, in which the initial set of image objects or ‘concepts’ is expanded using an external knowledge graph for generating diverse and informative stories. Despite these task-specific steps, we found that the vision encoder module in the architectures proposed for the various language generation tasks dealing with visual events share a common set of components that are broadly outlined in Figure 2.
@@
Figure 2: Outline of the architecture common across modeling approaches for tasks involving language generation from visual events.
2.2 The Vision-to-Language Bridge
@@
Some vision-to-language model architectures use an intermediate module that bridges the input and output modalities for effectively conditioning text generation on the extracted visual features. Different models operationalize this module with different degrees of complexity. Earlier approaches condition the text generation process by directly fusing vision encoder outputs with the language decoder input Kim et al. (2018). Some architectures employ cross-attention mechanisms to focus on the relevant parts of the visual features at various temporal positions during decoding Yao et al. (2015). However, approaches that adopt pre-trained models—e.g., CLIP-ViT-L Radford et al. (2021) as the visual model—tend to employ learnable intermediate layers for aligning and converting outputs of the vision encoder into a format that the language decoder can understand.
@@
In some of the proposed models, this intermediate module is a single linear layer that transforms the visual features into a common shared space, which can then be used by the language decoder Ko et al. (2023); Liu et al. (2023). In other models, advanced transformer-based projectors such as a Q-Former Li et al. (2023b) are used based on their ability to leverage cross-modal interactions effectively Han et al. (2024). In essence, Q-Former uses dynamic query vectors that are pre-trained to attend to both visual and textual representations, enhancing its ability to generalize and perform well (relative to a single linear layer) across different tasks. Besides these popular methods for adapting multimodal information, some approaches make use of graph neural networks for capturing relationships between objects in the images at different temporal positions and words in the corresponding sentences of the text Zhang and Peng (2019).
@@
While there is no definitive way to design this intermediate module, recent work has compared the two approaches, i.e., using cross-attention between modalities or using a multimodal projector for transforming vision encoder features into the language space, and found that the latter leads to improvements in performance of models Laurençon et al. (2024). This suggests that using intermediate multimodal projectors can help in extracting meaningful visual representations.
2.3 Language Decoder
@@
After encoding and adapting the visual information, models employ a language decoder component for text generation. The decoder can either be learned from scratch or consist of a pre-trained language model with additional trainable task-specific layers. Figure 2 summarizes the different ways in which this step is operationalized across tasks in the proposed architectures. Earlier models learn an RNN by initializing it with the visual context embedding from the previous steps Yao et al. (2015); Kim et al. (2018). The decoder then typically follows a ‘teacher forcing’ strategy during training to generate one word at a time autoregressively.
@@
Instead of training the decoder from scratch, some approaches use language models such as GPT-2 Radford et al. (2019) and LLAMA 2 Touvron et al. (2023), which are pre-trained on several text-only tasks such as question-answering and text classification. The pre-trained language models are either used directly for generation by freezing their parameters Han et al. (2023b, a, 2024), or by inserting and fine-tuning additional adaptive layers on top of them for ensuring relevance of the generated text to the downstream task of interest Yu et al. (2021). We also note that some models incorporate information from external knowledge bases into the decoder module to improve coherence and factuality of the generated text, e.g., TextKG Gu et al. (2023b) for the Video Captioning task and KG Story Hsu et al. (2020) for the Visual Storytelling task. This approach is shown to improve models in terms of correctness of answers, relevance of descriptions, and expressiveness of generated narratives. On the other hand, methods relying on external knowledge face several practical limitations, such as retrieval latency and model scalability. Furthermore, for applications in some domains (e.g., medical, financial), interacting with external sources may not be feasible due to privacy or security constraints.
@@
CARD Tu et al. (2024) ResNet, Transformer NA Transformer
SCORER+CBR Tu et al. (2023b) ResNet, MH(S/X)A NA Transformer
VARD-Trans Tu et al. (2023a) ResNet, Linear(s) NA Transformer
@@
ViLA Wang et al. (2025) ViT, Transformer Q-Former Flan-T5 XL
LLaMA-VQA Ko et al. (2023) CLIP-ViT-L Linear LLAMA
SeViLA Yu et al. (2023) ViT Q-Former Flan-T5 XL
FrozenBiLM Yang et al. (2022) CLIP-ViT-L Linear DeBERTa-V2-XL
@@
Vid2Seq Yang et al. (2023) CLIP ViT-L, Transformer NA T5-base
TextKG Gu et al. (2023b) Transformer NA Transformer
VTAR Shi et al. (2023) InceptionResNetV2, C3D Transformer Transformer
ENC-DEC Yao et al. (2015) 3DCNN Attention RNN
@@
MM-Narrator Zhang et al. (2024a) CLIP-ViT-L NA GPT-4
AutoAD-III Han et al. (2024) EVA-CLIP Q-Former LLAMA 2
AutoAD-II Han et al. (2023a) CLIP-ViT-L NA GPT-2, MHXA
AutoAD Han et al. (2023b) CLIP-ViT-L Transformer GPT-2
@@
MCSM+BART Chen et al. (2021) ResNet, RNN NA BART
TAPM Yu et al. (2021) ResNet, FRCNN NA GPT-2
KG Story Hsu et al. (2020) FRCNN NA Transformer
GLAC Net Kim et al. (2018) ResNet, RNN NA RNN
Table 1: A selection of recent models proposed for the tasks involving visual events considered in this paper. We only report models that are end-to-end and task-specific, i.e., trained or fine-tuned for the task at hand. For each model, we report the underlying vision encoder and language decoder, as well as the projector module, when applicable (NA: Not Applicable). Components with have been trained from scratch using only the datasets available for the corresponding task.
2.4 Off-the-shelf Pre-trained VLMs
The standard model architecture we have discussed so far is also generally present in powerful general-purpose foundation VLMs, which are pre-trained on several tasks using large amounts of data. Being multi-purpose by design, these models can be used directly for language generation tasks dealing with visual events. Their pre-training process typically happens in two stages—self-supervised alignment training and visual instruction tuning. During the first stage, only the parameters of the intermediate module connecting both unimodal backbones are updated (commonly using paired image-text data), using a contrastive training objective. In the second stage, models are instruction-tuned using multi-turn conversations obtained for visual data either through crowd-sourcing or by leveraging models such as GPT-4 OpenAI (2023).
Contrary to task-specific modeling approaches, these pre-trained VLMs are simply prompted (typically in a zero-shot manner) using visual tokens accompanied by task-specific instructions. Recent work has tested the ability of some of these pre-trained models to generate language from visual events Lin et al. (2024); Ye et al. (2025); Xie et al. (2024); Surikuchi, Fernández, and Pezzelle (2024). These multi-purpose models have been shown to be on par with or in some cases even outperform models that are designed and trained specifically for those tasks. Table 2 outlines the key components of some pre-trained VLMs used in a zero-shot manner for different vision-to-language tasks dealing with visual events.

### v3

# Natural Language Generation from Visual Events: State-of-the-Art and Key Open Questions

@@
Abstract.
In recent years, a substantial body of work in visually grounded natural language processing has focused on real-life multimodal scenarios such as describing content depicted in images or videos. However, comparatively less attention has been devoted to study the nature and degree of interaction between the different modalities in these scenarios. In this paper, we argue that any task dealing with natural language generation from sequences of images or frames is an instance of the broader, more general problem of modeling the intricate relationships between visual events unfolding over time and the features of the language used to interpret, describe, or narrate them. Therefore, solving these tasks requires models to be capable of identifying and managing such intricacies. We consider five seemingly different tasks, which we argue are compelling instances of this broader multimodal problem. Subsequently, we survey the modeling and evaluation approaches adopted for these tasks in recent years and examine the common set of challenges these tasks pose. Building on this perspective, we identify key open questions and propose several research directions for future investigation.
vision-to-language, language generation, visual sequences, visual storytelling, change captioning, movie auto audio description, video captioning, video question answering, multimodal NLP, visual events
$c$s: Computing methodologies Natural language generationccs: Computing methodologies Computer vision
1. Introduction
@@
Initially, much of the progress in visually grounded natural language processing (NLP) has focused on language generation from a static image. Tasks such as image captioning (Vinyals et al., 2015) and visual question answering (Antol et al., 2015) have demonstrated the potential of multimodal models to ground linguistic expressions in the visual modality (Baroni, 2016; Beinborn et al., 2018; Wang et al., 2024). However, the world is not static. Real-world applications such as media content creation, real-time sports commentary and news reporting, and interactive multimodal assistants frequently require reasoning about events as they unfold over time. This requires capturing key aspects of event understanding, including temporal dynamics, causality, and narrative structure, that arise from complex multimodal interaction.
@@
In this paper, we argue for a unified perspective: that all NLP tasks dealing with visual events, whether presented through videos or ordered image sets, are instances of a broader, more general problem. At the heart of this problem lies the need to model the intricate relationship between temporally ordered visual events and the structure, content, and function of the language used to interpret, describe, or narrate them. We claim that solving these tasks effectively requires models to be capable of identifying and managing such intricacies, leveraging the temporal, causal, and discourse-level patterns inherent in both visual data and language, as well as the many ways in which these aspects interact.
@@
To ground our argument, we examine five seemingly distinct tasks, each of which we claim exemplifies this broader multimodal challenge. Our aim is to provide a survey of modeling approaches and evaluations proposed for a few, selected relevant tasks and to illustrate the general problem of generating language from visual events. The tasks we consider are Change Captioning (Jhamtani and Berg-Kirkpatrick, 2018), Video Question Answering (Video QA) (Zeng et al., 2017; Bansal et al., 2020), Video Captioning (Yao et al., 2015), Movie Auto Audio Description (Movie Auto AD) (Han et al., 2023b, a), and Visual Storytelling (Huang et al., 2016). Figure 1 presents an illustration and a prototypical example for each of the five tasks, drawn from publicly available datasets. By analyzing them through a shared lens, we highlight common modeling and evaluation challenges that are often overlooked when tasks are considered in isolation. Yet, we recognize that these tasks also differ in crucial ways that stem from their distinct communicative intents and objectives. For instance, crafting a coherent and imaginative narrative in Visual Storytelling poses fundamentally different demands compared to pinpointing specific visual changes in Change Captioning. These task-specific nuances, which can be ignored or flattened when focusing solely on shared technical challenges, underscore the need for specific modeling proposals and evaluation protocols, which is the second key argument of our paper.
@@
  • We offer a comprehensive survey of modeling approaches, training paradigms, and evaluation methodologies used across five vision-to-language tasks involving visual events. By identifying shared pillars and methodological commonalities, we provide a unified lens that we hope will serve as a foundation for future research.
@@
  • We show that, within the problem of generating language from visual events, different tasks vary along various features based on their communicative intents. To this end, we provide a quantitative characterization of the tasks based on two dimensions—visual similarity and textual consistency—capturing the similarity of the elements in the visual modality, and the consistency of the textual modality.
  • We outline several promising research directions aimed at advancing models’ ability to capture the complex, time-dependent relationships between visual and linguistic modalities. These include cognitively inspired representations, improved discourse modeling, and principled evaluation metrics that reflect narrative quality.
@@
¡Write a story for this image sequence.,
@@
On the left, this figure shows a sequence of images or video frames positioned side-by-side. On the right, text corresponding to the visual inputs is provided.
[…]
2. Model Architectures
Modeling approaches to text generation from visual events have evolved over time from being based on recurrent neural networks (RNN; Hochreiter and Schmidhuber, 1997) to being transformer-based (Vaswani et al., 2017). More recent models directly leverage pre-trained large (vision)-language models (LLMs/VLMs), often in a zero-shot manner. In this section, we discuss this evolution and summarize the various state-of-the-art model architectures proposed for this problem. We ground our discussion in five representative tasks—listed in Figure 1—that serve to exemplify the main trends. Change Captioning requires generating a caption to describe semantic changes between a pair of related input images. This task is particularly relevant for applications involving surveillance or quality inspection of ‘before’ and ‘after’ visual data. The goal of Visual Storytelling is to generate a fictitious, coherent narrative from a temporally-ordered sequence of images or frames. Applications of this task include digital storytelling and media creation. Video Captioning concerns describing videos by summarizing them, while Video QA involves generating appropriate answers to questions about the entities and events in videos. Both tasks have a wide range of applications, such as content summarization on streaming platforms and interacting with virtual assistants. Movie Auto AD involves generating narrative descriptions to complement the dialogues and soundtracks of movies. Enhancing the accessibility of movies and automatizing dubbing are some of the applications of this task.
@@ 2. Model Architectures
2.1. Vision Encoder
The primary purpose of a vision encoder in vision-to-language tasks is to extract information from the visual input and to optimally encode it into a contextual representation that guides language generation. To achieve this, encoders in the proposed models follow multiple steps, some of which are common across language generation tasks from visual events. First, a pre-trained vision model is utilized for extracting feature representations of the raw input sequences of images or video frames. Earlier approaches used vision models based on convolutional neural networks (CNNs) such as ResNet (He et al., 2016) or R3D (Tran et al., 2018) that are primarily pre-trained on the object detection task using large amounts of image or video data. Most of the recent models across tasks use transformer-based vision models pre-trained for various image-only and image-text alignment objectives, e.g., CLIP-ViT-L (Radford et al., 2021). We note that besides the primary input sequence of images or video frames, models proposed for some of the tasks, e.g., Movie Auto AD, utilize additional input data such as close-ups of characters in the movie clips (exemplars) (Han et al., 2023a). Furthermore, the TAPM (Yu et al., 2021) model proposed for the Visual Storytelling task utilizes FasterRCNN (Ren et al., 2015) for extracting such local entity-level features alongside the global image-level features from ResNet.
@@ 2.1. Vision Encoder
Beyond these common components across tasks, vision-encoders may also contain additional task-specific steps for capturing the visual information in a way that suits the task’s objective better. For instance, the ViLA model (Wang et al., 2025) for the Video QA task utilizes a learnable Frame-Sampler component to efficiently select a small subset of frames that are most likely to contain the relevant information needed to answer the question. Another example with a task-specific step is the MSCM+BART model (Hsu et al., 2020) for Visual Storytelling, in which the initial set of image objects or ‘concepts’ is expanded using an external knowledge graph for generating diverse and informative stories. Despite these task-specific steps, we found that the vision encoder module in the architectures proposed for the various language generation tasks dealing with visual events share a common set of components that are broadly outlined in Figure 2.
[…]
@@
Figure 2. Outline of the architecture common across modeling approaches for tasks involving language generation from visual events.
2.2. The Vision-to-Language Bridge
@@
Some vision-to-language model architectures use an intermediate module that bridges the input and output modalities for effectively conditioning text generation on the extracted visual features. Different models operationalize this module with different degrees of complexity. Earlier approaches condition the text generation process by directly fusing vision encoder outputs with the language decoder input (Kim et al., 2018). Some architectures employ cross-attention mechanisms to focus on the relevant parts of the visual features at various temporal positions during decoding (Yao et al., 2015). However, approaches that adopt pre-trained models—e.g., CLIP-ViT-L (Radford et al., 2021) as the visual model—tend to employ learnable intermediate layers for aligning and converting outputs of the vision encoder into a format that the language decoder can understand.
@@
In some of the proposed models, this intermediate module is a single linear layer that transforms the visual features into a common shared space, which can then be used by the language decoder (Ko et al., 2023; Liu et al., 2023). In other models, advanced transformer-based projectors such as a Q-Former (Li et al., 2023a) are used based on their ability to leverage cross-modal interactions effectively (Han et al., 2024). In essence, Q-Former uses dynamic query vectors that are pre-trained to attend to both visual and textual representations, enhancing its ability to generalize and perform well (relative to a single linear layer) across different tasks. Besides these popular methods for adapting multimodal information, some approaches make use of graph neural networks for capturing relationships between objects in the images at different temporal positions and words in the corresponding sentences of the text (Zhang and Peng, 2019).
@@
While there is no definitive way to design this intermediate module, recent work has compared the two approaches, i.e., using cross-attention between modalities or using a multimodal projector for transforming vision encoder features into the language space, and found that the latter leads to improvements in performance of models (Laurençon et al., 2024). This suggests that using intermediate multimodal projectors can help in extracting meaningful visual representations.
2.3. Language Decoder
@@
After encoding and adapting the visual information, models employ a language decoder component for text generation. The decoder can either be learned from scratch or consist of a pre-trained language model with additional trainable task-specific layers. Figure 2 summarizes the different ways in which this step is operationalized across tasks in the proposed architectures. Earlier models learn an RNN by initializing it with the visual context embedding from the previous steps (Yao et al., 2015; Kim et al., 2018). The decoder then typically follows a ‘teacher forcing’ strategy during training to generate one word at a time autoregressively.
@@
Instead of training the decoder from scratch, some approaches use language models such as GPT-2 (Radford et al., 2019) and LLAMA 2 (Touvron et al., 2023), which are pre-trained on several text-only tasks such as question-answering and text classification. The pre-trained language models are either used directly for generation by freezing their parameters (Han et al., 2023b, a, 2024), or by inserting and fine-tuning additional adaptive layers on top of them for ensuring relevance of the generated text to the downstream task of interest (Yu et al., 2021). We also note that some models incorporate information from external knowledge bases into the decoder module to improve coherence and factuality of the generated text, e.g., TextKG (Gu et al., 2023a) for the Video Captioning task and KG Story (Hsu et al., 2020) for the Visual Storytelling task. This approach is shown to improve models in terms of correctness of answers, relevance of descriptions, and expressiveness of generated narratives. On the other hand, methods relying on external knowledge face several practical limitations, such as retrieval latency and model scalability. Furthermore, for applications in some domains (e.g., medical, financial), interacting with external sources may not be feasible due to privacy or security constraints.
@@
CARD (Tu et al., 2024) ResNet, Transformer NA Transformer
SCORER+CBR (Tu et al., 2023b) ResNet, MH(S/X)A NA Transformer
VARD-Trans (Tu et al., 2023a) ResNet, Linear(s) NA Transformer
@@
ViLA (Wang et al., 2025) ViT, Transformer Q-Former Flan-T5 XL
LLaMA-VQA (Ko et al., 2023) CLIP-ViT-L Linear LLAMA
SeViLA (Yu et al., 2023) ViT Q-Former Flan-T5 XL
FrozenBiLM (Yang et al., 2022) CLIP-ViT-L Linear DeBERTa-V2-XL
@@
Vid2Seq (Yang et al., 2023) CLIP ViT-L, Transformer NA T5-base
TextKG (Gu et al., 2023a) Transformer NA Transformer
VTAR (Shi et al., 2023) InceptionResNetV2, C3D Transformer Transformer
ENC-DEC (Yao et al., 2015) 3DCNN Attention RNN
@@
MM-Narrator (Zhang et al., 2024a) CLIP-ViT-L NA GPT-4
AutoAD-III (Han et al., 2024) EVA-CLIP Q-Former LLAMA 2
AutoAD-II (Han et al., 2023a) CLIP-ViT-L NA GPT-2, MHXA
AutoAD (Han et al., 2023b) CLIP-ViT-L Transformer GPT-2
@@
MCSM+BART (Chen et al., 2021) ResNet, RNN NA BART
TAPM (Yu et al., 2021) ResNet, FRCNN NA GPT-2
KG Story (Hsu et al., 2020) FRCNN NA Transformer
GLAC Net (Kim et al., 2018) ResNet, RNN NA RNN
Table 1. A selection of recent models proposed for the tasks involving visual events considered in this paper. We only report models that are end-to-end and task-specific, i.e., trained or fine-tuned for the task at hand. For each model, we report the underlying vision encoder and language decoder, as well as the projector module, when applicable (NA: Not Applicable). Components with have been trained from scratch using only the datasets available for the corresponding task.
2.4. Off-the-shelf Pre-trained VLMs
The standard model architecture we have discussed so far is also generally present in powerful general-purpose foundation VLMs, which are pre-trained on several tasks using large amounts of data. Being multi-purpose by design, these models can be used directly for language generation tasks dealing with visual events. Their pre-training process typically happens in two stages—self-supervised alignment training and visual instruction tuning. During the first stage, only the parameters of the intermediate module connecting both unimodal backbones are updated (commonly using paired image-text data), using a contrastive training objective. In the second stage, models are instruction-tuned using multi-turn conversations obtained for visual data either through crowd-sourcing or by leveraging models such as GPT-4 (OpenAI, 2023).
Contrary to task-specific modeling approaches, these pre-trained VLMs are simply prompted (typically in a zero-shot manner) using visual tokens accompanied by task-specific instructions. Recent work has tested the ability of some of these pre-trained models to generate language from visual events (Lin et al., 2024; Ye et al., 2025; Xie et al., 2024; Surikuchi et al., 2024). These multi-purpose models have been shown to be on par with or in some cases even outperform models that are designed and trained specifically for those tasks. Table 2 outlines the key components of some pre-trained VLMs used in a zero-shot manner for different vision-to-language tasks dealing with visual events.